Find the point on the y-axis that is equidistant from (5,1) and (-3,-1).
Answer by stanbon(75887)

Find the point on the y-axis that is equidistant from (5,1) and (-3,-1).
The x and the y coordinate of the midpoint are the average of the x and the
y coordinates of the end points.
------------
x= (5-3)/2 = 1
y= (1-1)/2 = 0
-------------------
Midpoint : (1,0)
=================
Cheers,
Stan H.